# Large Language Models and Impossible Language Acquisition: “False Promise” or an Overturn of our Current Perspective towards AI

### v1

@@ Abstract
In Chomsky’s provocative critique “The False Promise of CHATGPT,” Large Language Models (LLMs) are characterized as mere pattern predictors that do not acquire languages via intrinsic causal and self-correction structures like humans, therefore are not able to distinguish impossible languages. It stands as a representative in a fundamental challenge to the intellectual foundations of AI, for it integrally synthesizes major issues in methodologies within LLMs and possesses an iconic a priori rationalist perspective. We examine this famous critic from both the perspective in pre-existing literature of linguistics and psychology as well as a research based on an experiment inquiring the capacity of learning both possible and impossible languages among LLMs. We constructed a set of syntactically impossible languages by applying certain transformations to English. These include reversing whole sentences, and adding negation based on word-count parity. Two rounds of controlled experiments were each conducted on GPT-2 small models and long short-term memory (LSTM) models. Statistical analysis (Welch’s t-test) shows GPT2 small models underperform in learning all of the impossible languages compared to their performance on the possible language ($p<.001$). On the other hand, LSTM models’ performance tallies with Chomsky’s argument, suggesting the irreplaceable role of the evolution of transformer architecture. Based on theoretical analysis and empirical findings, we propose a new vision within Chomsky’s theory towards LLMs, and a shift of theoretical paradigm outside Chomsky, from his “rationalist-romantics” paradigm to functionalism and empiricism in LLMs research.
@@ 1.2.1 In Theoretical Aspect
In this fundamental discussion, we engage with classic literature from both developmental psychology and philosophy, which both offer alternative understandings of Chomsky’s questions, furthermore, new theoretical architectures. For instance, we have drawn upon Jean Piaget’s constructivist theory of cognitive development (Piaget, 1970), which describes the staging progress of human cognition, engaging with the adaptation to unfamiliar experience.
@@
Although being widely accepted, the PoS Argument (PoSA) that props the Chomskyan UG is not flawless even in pure reasoning. Recent literature review straightforwardly challenges the establishment of PoSA, and displays the dilemma for it to back up linguistic innatism (Skidelsky, 2016). Whereas studies concerning the learnability of language in statistical learning (Lewis and Elman, 2001; Perfors et al., 2011) have shown AIs’ real performances in language acquisition as learners with similarities to human, rather than the prima facie that regard them as parroting programs.
@@ 1.2.3 A Clarification of our Intention
Although it appears that we are standing on the opposite side to the Chomskyan school, we are not supporting an environmentalism based on the Anglo-American empiricist tradition, which believes that “the mind results from a few simple operations of association on the basis of contiguity,” and for humans alone any intellect barrier associated with language acquisition can be overcome by learning in whatever means (Chomsky, 1973). We are not contesting theories on the unique biological structures that inherently empower humans to acquire languages. In Chomsky (2005), the biological view on environment and experience could represent a consensus between us and Chomsky. Though we suggest that there is comparability between the learning process in humans and in LLMs, the essential differences between them apparently state the impossibility of transplanting every conclusion from machine learning to humans, regardless of the external survey on LLMs. This study only intends to focus on the problems of Chomsky’s critiques within practical considerations for the future development of machine learning, and what his limitations have revealed in his cognitive theory.
@@ 1.2.4 Preparation and Methodology of the Experiment
The aforementioned theoretical juxtaposition still requires fundamental evidence to internally reconsider Chomsky’s criticism that can be empirically measured. Such supporting evidence is already concluded by a few pre-existing researches, which show the universality of related experience, enhancing its influence on the general methodological choices. This evidence is the underperformance of LLMs when learning impossible languages. In Chomsky’s original critique, the core argument (a) is directly relevant to the issue. It is claimed that LLMs are merely pattern predictors, unable to distinguish an impossible language, but, on the contrary, both the experiments that we performed and those from Kallini et al. (2024) suggest the hardship of LLMs when acquiring impossible languages compared to their capacity in learning natural languages.
The experiment that takes place in our paper also intends to examine the claim that LLMs are equally adequate to learn both possible and impossible languages, judged by the Chomskyan studies on syntactic structures. Kallini et al. (2024) have provided a mature basic design to evaluate and quantify the performance of LLMs in natural languages and impossible languages. Our experiment is partly based on that method, but also based on our original improvements: firstly, we chose not only the BabyLM dataset (Warstadt et al., 2023) as the corpus for our training, for it is not always pragmatic to implement impossible grammatical changes on complex sentences while leaving the linguistically unnatural characteristics apparent for LLMs. It is essential to avoid difficulties such as word segmentation, long-sentence structures, and other factors that could disturb LLMs’ acquisition process, and exclude other variables other than the possibility of each language itself. Thus, we used an automatic mechanism to generate simple SVO sentences for every grammatical transformation, which is also beneficial to strengthen the universality and validity of our findings in a larger variety of datasets other than BabyLM, as well as in making it easier for researchers who lack hardware resources like ourselves. Secondly, in the precise choices of indicators for the effectiveness of learning processes, we applied the data collection of both perplexity and loss value, whereas Kallini et al. (2024) only focused on perplexity. In this way, our experiment can be more objective in the deduction of conclusions.
@@ 1.2.5 Our Empirical Assessment
At first, our experiment was designed to verify Chomsky’s opinion on LLMs. The primal hypothesis was that the models of each kind trained on both possible and impossible languages can achieve the same average perplexity, proving that LLMs are incapable of identifying natural and unnatural languages. However, our overall experimental results ended up suggesting that the control group has achieved the lowest loss value throughout the whole training, while its perplexity was also one of the lowermost. Two rounds of experiments conducted on two large datasets have concluded the same result: GPT-2 models are biased to possible languages. They are more efficient to learn possible languages compared to their underperformance when acquiring impossible languages ($p<.001$). Same experiments were again conducted on LSTM models, where their performances from each group are similar to each other. [^1] The code for the experiment can be found at https://github.com/xieguaiwu/LLMs-and-impossible-language-acquisition
@@ 2.1.1 Line of Reasoning
Undeniably, language competence must have a physiological and physical foundation. However, limited empirical knowledge is not enough evidence to consequentially describe this foundation as “a ‘language acquisition device’ that takes experience as ‘input’ and gives the language as an ‘output’.”’ (Chomsky, 2000) Apparently, any insight we can have on the mind is bounded on observable phenomenon, from nervous activities to verbal communications. Numerous mind philosophers have proposed divergent theories with different tendencies as well as searchable logical discontinuities, due to a contradiction between the desire for a universal description, and the chaotic reality that leaves little possibility to experientially find one, where nothing can make one theory truer than the others. When logical discontinuities in a seemingly complete theory such as the one from Chomsky are found, it is likely that those discontinuities are arise from the conceptualization of observations, which itself structuralizes more that can be supported by the evidence.
In Chomsky’s example, he structuralizes a pure and absolute metaphysical innateness from early children verbal behavior (Chomsky, 1973, 1988, 2000, 2002, 2009), and then reinforces his criticism on LLMs with such conceptualization that lacks sufficient empirical support in the context of machine learning, despite that there are also direct observations that can support alternative views as well, which would be an intuitive and experiential support without enough statistical effectiveness, just like that in Chomsky works. Same insight on Chomsky’s methodology can also be seen in Hintikka (1999), where it was regarded as the origin of the use of intuitions in philosophical argumentation after the mid-1960s.
@@ 2.1.2 External Analysis
Philosophically, the problem with Chomsky’s argument is similar to what Quine criticizes logical positivism. Quine targets on the insufficient distinction between analytic and synthetic statement, especially the one drawn by Carnap’s artificial language, which is: “a distinction to be drawn at all is an unempirical dogma of empiricists, a metaphysical article of faith.” (Quine, 1951) In the same meaning, Chomsky also draws such a line in the middle of language faculty and language performance, where he thereby divides his theory into two parts – language competence and syntactic analysis. In reality, any hint of a language faculty that is stored inside a “language organ” is given by language performance.
@@
A classic study also supports the learnability of language in statistical learning. It was found in Lewis and Elman (2001) that even simple recurrent networks (SRNs) can capture the inner structure of relative clauses, suggesting the emergence of syntactic structural understanding. On the same level, Perfors et al. (2011) indicated with Bayesian architecture that a learner with only general inductive is already able to identify hierarchical phrase structures. Furthermore, they clarified the problem of PoSA in practice: what should be focus on is “what kind of knowledge must be assumed as an innate constraint on the learner’s inductive hypotheses, rather than on what kind of representational machinery must be available to the learner.” (Perfors et al., 2011)
@@ 2.3.1 Impossible Language Training
Moro (2016) and Moro et al. (2023) defines impossible language as non-hierarchical and linear (see Figure 2), radically divergent from the recursive property in Chomsky’s description (Chomsky, 1957, 1965). In the context-free grammar (CFG) of Syntactic Structures particularly, recursiveness is an essential tool in order to proceed a derivation using limited rewrite rules. For instance, in a common CFG denoted as
@@
(1) The horse has enjoyed the school ($T+N+S+Aux+V+T+N+\emptyset$)
(1) NOT The horse has enjoyed the school
(2) The girl is given cats ($T+N+S+Aux+V+N+S$)
(2) The girl is given cats NOT
@@ 3 Towards a Functional Perspective
According to our experiment, LLMs are not intrinsically unbiased in language learning, like what is depicted by Chomsky’s original critique; while in the broader cognitive practice of machine learning, the learning mechanism of each architecture have great impact on the actual behavior of language acquisition. The leap from RNNs to the transformer architecture has influenced their performances in impossible language learning, which is indispensable for the revolutionary achievement. Though a scientific breakthrough would call for an internal exploration, but to what extend can we take our understanding on the mechanism of intellect as the intellect itself, assume that we could take this step in the future?
@@
In an “expression-wielding process,” the subject is given rules to follow, lacking the intellect of generalization. Both Chomsky and Ryle point this power as creativity, where the former posits it within its philosophical innateness, arguing it mysteriously as something only humans can possess, without empirical evidence. Such doctrine from Cartesian is a belief, instead of effective testimony that can support a total denial on LLMs. Getting rid of the ambiguity caused by metaphysical arguments in scientific investigations is needful, while keep complicating the discourse is unnecessary. What the future of AI needs is a philosophical foundation for scientific methodology, which studies LLMs as a cognitive subject. Under this consideration, Ryle’s logical behaviorism shares a common ground with scientific approach, for it regards the mind as an experimental subject, and because it bounds the role of philosophy as nothing but a clarification, leaving the leadership of theoretical investigation to empirical evidence.
Wei et al. (2024) is such supporting evidence within Ryle’s work. The introduction of chain-of-thought prompting is proven to be significantly helpful in multi-step reasoning resembling procedural knowledge application. As an indicator of the construction of theories, the empirical conclusions are able to question opinions that claim LLMs as static knowledge storages. Furthermore, in Guo et al. (2025), the pure reinforcement learning (RL) of DeepSeek-R1 model directly embodies a Piagetian paradigm of cognitive development. The most notable clue of creativity and adaptation is reflected in the observed reasoning strategies, which are self-evolved via the learning mechanism itself. Such dynamic interactions are emerged, beyond the Chomskyan theoretical limitation.
In conclusion, the functional perspective combines the epistemology of Piaget and Ryle, which conducts a shift from Chomsky’s “rationalist-romantics” paradigm. The shift represents an evolution in the measurement of a cognitive system, challenging the traditional nativist understanding, but advocates the indicators like efficacy, and sophistication of for a cognitive subject in specific problem-oriented actions.
@@ 4.1 On GPT-2 Small Models as Experimental Subjects
Apart from questions towards the technical details of impossible languages, it is also important to explain the use of GPT-2 small models, instead of stronger models. We believe that GPT-2 small model plays an essential role in language acquisition experiments of machine learning as a guinea pig or white mice in biological ones.
@@
In general, studying language learning on GPT-2 is analogous to experiments on white mouse to test vaccines. Both of which utilize the controllability in puny experimental subjects, and generalize findings from a miniature test to a larger scale. This operation represents the introduction of scientific and empirical methodology in AI.

### v4

@@ Abstract
In Chomsky’s provocative critique “The False Promise of CHATGPT,” Large Language Models (LLMs) are characterized as mere pattern predictors that do not acquire languages via intrinsic causal and self-correction structures like humans, therefore are not able to distinguish impossible languages. It stands as a representative in a fundamental challenge to the intellectual foundations of AI, for it integrally synthesizes major issues in methodologies within LLMs and possesses an iconic a priori rationalist perspective. We examine this famous critique from both the perspective in pre-existing literature of linguistics and psychology as well as a research based on an experiment inquiring into the capacity of learning both possible and impossible languages among LLMs. We constructed a set of syntactically impossible languages by applying certain transformations to English. These include reversing whole sentences, and adding negation based on word-count parity. Two rounds of controlled experiments were each conducted on GPT-2 small models and long short-term memory (LSTM) models. Statistical analysis (Welch’s t-test) shows GPT2 small models underperform in learning all of the impossible languages compared to their performance on the possible language ($p<.001$). On the other hand, LSTM models’ performance tallies with Chomsky’s argument, suggesting the irreplaceable role of the evolution of transformer architecture. Based on theoretical analysis and empirical findings, we propose a new vision within Chomsky’s theory towards LLMs, and a shift of theoretical paradigm outside Chomsky, from his “rationalist-romantics” paradigm to functionalism and empiricism in LLMs research.
@@ 1.2.1 In Theoretical Aspect
In this fundamental discussion, we engage with classic literature from both developmental psychology and philosophy, which both offer alternative understandings of Chomsky’s questions, and furthermore, new theoretical architectures. For instance, we have drawn upon Jean Piaget’s constructivist theory of cognitive development (Piaget, 1970), which describes the staging progress of human cognition, engaging with the adaptation to unfamiliar experience.
@@
Although being widely accepted, the PoS Argument (PoSA) that props the Chomskyan UG is not flawless even in pure reasoning. A recent literature review challenges the establishment of PoSA, and displays the dilemma for it to back up linguistic innatism (Skidelsky, 2016). Whereas studies concerning the learnability of language in statistical learning (Lewis and Elman, 2001; Perfors et al., 2011) have shown AIs’ real performance in language acquisition as learners with similarities to human, rather than the prima facie that regard them as parroting programs.
@@ 1.2.3 A Clarification of our Intention
Although it appears that we are standing on the opposite side from the Chomskyan school, we are not supporting an environmentalism based on the Anglo-American empiricist tradition, which believes that “the mind results from a few simple operations of association on the basis of contiguity,” and for humans alone any intellectual barrier associated with language acquisition can be overcome by learning by whatever means (Chomsky, 1973). We are not contesting theories of the unique biological structures that inherently empower humans to acquire languages. The biological view on environment and experience in Chomsky (2005) could represent a consensus between us and Chomsky. Though we suggest that there is comparability between the learning process in humans and in LLMs, the essential differences between them apparently demonstrate the impossibility of transplanting every conclusion from machine learning to humans, regardless of the external survey on LLMs. The only kinds of comparability between humans and LLMs we can accept are those shown in intellectual behaviors, which do not contrast with their neurological incommensurability, for LLMs possess no biological substrate, but hold comprehensive experience. This study only intends to focus on the problems of Chomsky’s critiques within practical considerations for the future development of machine learning, and what his limitations have revealed in his cognitive theory.
@@ 1.2.4 Preparation and Methodology of the Experiment
The aforementioned theoretical juxtaposition still requires fundamental evidence to internally reconsider Chomsky’s criticism that can be empirically measured. Such supporting evidence is already concluded by a few pre-existing studies, which show the universality of related experience, enhancing its influence on the general methodological choices. This evidence is the underperformance of LLMs when learning impossible languages. In Chomsky’s original critique, the core argument (a) is directly relevant to the issue. It is claimed that LLMs are merely pattern predictors, unable to distinguish an impossible language, but, on the contrary, both the experiments that we performed and those from Kallini et al. (2024) suggest the hardship of LLMs when acquiring impossible languages compared to their capacity in learning natural languages.
The experiment that takes place in our paper also intends to examine the claim that LLMs are equally adequate to learn both possible and impossible languages, judged by the Chomskyan studies on syntactic structures. Kallini et al. (2024) have provided a mature basic design to evaluate and quantify the performance of LLMs in natural languages and impossible languages. Our experiment is partly based on that method, but also based on our original improvements: firstly, we chose not only the BabyLM dataset (Warstadt et al., 2023) as the corpus for our training, for it is not always practical to implement impossible grammatical changes on complex sentences while leaving the linguistically unnatural characteristics apparent for LLMs. It is essential to avoid difficulties such as word segmentation, long-sentence structures, and other factors that could disturb LLMs’ acquisition process, and exclude variables other than the possibility of each language itself. Thus, we used an automatic mechanism to generate simple SVO sentences for every grammatical transformation, which is also beneficial to strengthen the universality and validity of our findings in a larger variety of datasets other than BabyLM, as well as in making it easier for researchers who lack hardware resources like ourselves. Secondly, in the precise choices of indicators for the effectiveness of learning processes, we applied the data collection of both perplexity and loss value, whereas Kallini et al. (2024) only focused on perplexity. In this way, our experiment can be more objective in the deduction of conclusions.
@@ 1.2.5 Our Empirical Assessment
At first, our experiment was designed to verify Chomsky’s opinion on LLMs. The primal hypothesis was that the models of each kind trained on both possible and impossible languages can achieve the same average perplexity, proving that LLMs are incapable of identifying natural and unnatural languages. However, our overall experimental results ended up suggesting that the control group has achieved the lowest loss value throughout the whole training, while its perplexity was also one of the lowermost. Two rounds of experiments conducted on two large datasets have yielded the same result: GPT-2 models are biased to possible languages. They are more efficient to learn possible languages compared to their underperformance when acquiring impossible languages ($p<.001$). The same experiments were again conducted on LSTM models, where their performances from each group are similar to each other. [^1] The code for the experiment can be found at https://github.com/xieguaiwu/LLMs-and-impossible-language-acquisition
@@ 2.1.1 Line of Reasoning
Undeniably, language competence must have a physiological and physical foundation. However, limited empirical knowledge is not enough evidence to consequentially describe this foundation as “a ‘language acquisition device’ that takes experience as ‘input’ and gives the language as an ‘output’.”’ (Chomsky, 2000) Apparently, any insight we can have on the mind is bounded on observable phenomenon, from nervous activities to verbal communications. Numerous mind philosophers have proposed divergent theories with different tendencies as well as searchable logical discontinuities, due to a contradiction between the desire for a universal description, and the chaotic reality that leaves little possibility to experientially find one, where nothing can make one theory truer than the others. When logical discontinuities in a seemingly complete theory such as the one from Chomsky are found, it is likely that those discontinuities arise from the conceptualization of observations, which itself structuralizes more that can be supported by the evidence.
In Chomsky’s example, he structuralizes a pure and absolute metaphysical innateness from early children’s verbal behavior (Chomsky, 1973, 1988, 2000, 2002, 2009), and then reinforces his criticism on LLMs with such conceptualization that lacks sufficient empirical support in the context of machine learning, despite that there are also direct observations that can support alternative views as well, which would be an intuitive and experiential support without enough statistical effectiveness, just like that in Chomsky works. Same insight on Chomsky’s methodology can also be seen in Hintikka (1999), where it was regarded as the origin of the use of intuitions in philosophical argumentation after the mid-1960s.
@@ 2.1.2 External Analysis
Philosophically, the problem with Chomsky’s argument is similar to what Quine criticizes logical positivism. Quine targets the insufficient distinction between analytic and synthetic statement, especially the one drawn by Carnap’s artificial language, which is: “a distinction to be drawn at all is an unempirical dogma of empiricists, a metaphysical article of faith.” (Quine, 1951) In the same meaning, Chomsky also draws such a line in the middle of language faculty and language performance, where he thereby divides his theory into two parts – language competence and syntactic analysis. In reality, any hint of a language faculty that is stored inside a “language organ” is given by language performance.
@@
A classic study also supports the learnability of language in statistical learning. It was found in Lewis and Elman (2001) that even simple recurrent networks (SRNs) can capture the inner structure of relative clauses, suggesting the emergence of syntactic structural understanding. On the same level, Perfors et al. (2011) indicated with Bayesian architecture that a learner with only general inductive abilities is already able to identify hierarchical phrase structures. Furthermore, they clarified the problem of PoSA in practice: what should be focused on is “what kind of knowledge must be assumed as an innate constraint on the learner’s inductive hypotheses, rather than on what kind of representational machinery must be available to the learner.” (Perfors et al., 2011)
@@ 2.2.3 Impossible Language Training
Moro (2016) and Moro et al. (2023) define impossible language as non-hierarchical and linear (see Figure 2), radically divergent from the recursive property in Chomsky’s description (Chomsky, 1957, 1965). Moreover, Musso et al. (2003) points to the existence of a clinical difference for humans to learn possible and impossible languages, suggesting that Broca’s area, an essential part of the brain to fulfill speech production functions, “is progressively disengaged” when dealing with unreal grammars. Andrea Moro’s other important works are also dedicated to discussing the neurological foundation of impossible languages identification: in Moro et al. (2001), the tripartite structure of language (Chomsky, 1957, 1965, 2014) is discovered to have a corresponding distributed neural network implementation, which relates to the subjects’ performances in the exposure of “pseudosentences”. Such profound studies provide additional dimensions for the biological content of impossible languages; whereas Tettamanti et al. (2002) provides the first direct neuroimaging evidence, which identifies neural correlates for the acquisition of grammatical and nongrammatical rules in adults, and shows a higher competence in Broca’s area while handling grammatical rules. The specific similarities of impossible language acquisition between LLMs that are based on transformer architecture and humans could suggest a deep comparability in their linguistic learning process.
Back to a formal divergence of possible and impossible languages, in the context-free grammar (CFG) of Syntactic Structures particularly, recursiveness is an essential tool in order to proceed a derivation using limited rewrite rules. For instance, in a common CFG denoted as
@@
(1a) The horse has enjoyed the school ($T+N+S+Aux+V+T+N+\emptyset$)
(1b) NOT The horse has enjoyed the school
(2a) The girl is given cats ($T+N+S+Aux+V+N+S$)
(2b) The girl is given cats NOT
@@
2.2.4 Syntactic Surprisal and Hierarchical Bias
A recent study provides further support for the architectural significance in language acquisition, comparing the effects of four probability models based on surprisal, revealing that only surprisal models containing hierarchical structures and the categorization of words in part-of-speech classes that suggest the linguistic property of certain words through a rather semantically insensitive manner (Greco et al., 2024). Although the experiments conducted in that research choose n-gram models, their conclusion aligns with the discoveries from McCoy et al. (2020), and also our empirical observations. GPT2 small models with attention mechanisms that differ from LSTM models that can only operate on sequential processing also demonstrate a preference to hierarchical languages, which is additional evidence for the two remarkable previous works. Together, we can unanimously come to a synthetical understanding: it is not needful to draw a line between statistical learning and Chomskyan linguistics. Methodologically, this reduction of dogmatic contrast is similar to Quine (1951), as mentioned before in section 2.1.2. The dichotomy in Chomsky (2023) is challenged by such reconsideration.
@@ 3 Towards a Functional Perspective
According to our experiment, LLMs are not intrinsically unbiased in language learning, like what is depicted by Chomsky’s original critique; while in the broader cognitive practice of machine learning, the learning mechanism of each architecture has great impact on the actual behavior of language acquisition. The leap from RNNs to the transformer architecture has influenced their performances in impossible language learning, which is indispensable for the revolutionary achievement. Though a scientific breakthrough would call for an internal exploration, to what extent can we take our understanding on the mechanism of intellect as the intellect itself, assume that we could take this step in the future?
@@
In an “expression-wielding process,” the subject is given rules to follow, lacking the intellect of generalization. Both Chomsky and Ryle point this power as creativity, where the former posits it within its philosophical innateness, arguing it mysteriously as something only humans can possess, without empirical evidence. Such doctrine from Cartesian is a belief, instead of effective testimony that can support a total denial on LLMs. Getting rid of the ambiguity caused by metaphysical arguments in scientific investigations is needful, while keeping the discourse complicated is unnecessary. What the future of AI needs is a philosophical foundation for scientific methodology, which studies LLMs as a cognitive subject. Under this consideration, Ryle’s logical behaviorism shares a common ground with scientific approach, for it regards the mind as an experimental subject, and because it bounds the role of philosophy as nothing but a clarification, leaving the leadership of theoretical investigation to empirical evidence.
Wei et al. (2024) is such supporting evidence within Ryle’s work. The introduction of chain-of-thought prompting is proven to be significantly helpful in multi-step reasoning resembling procedural knowledge application. As an indicator of the construction of theories, the empirical conclusions are able to question opinions that claim LLMs as static knowledge storages. Furthermore, in Guo et al. (2025), the pure reinforcement learning (RL) of DeepSeek-R1 model directly embodies a Piagetian paradigm of cognitive development. The most notable clue of creativity and adaptation is reflected in the observed reasoning strategies, which are self-evolved via the learning mechanism itself. Such dynamic interactions emerge, beyond the Chomskyan theoretical limitation.
In conclusion, the functional perspective combines the epistemology of Piaget and Ryle, which conducts a shift from Chomsky’s “rationalist-romantics” paradigm. The shift represents an evolution in the measurement of a cognitive system, challenging the traditional nativist understanding, but advocates the indicators like efficacy, and sophistication of a cognitive subject in specific problem-oriented actions.
@@ 4.1 On GPT-2 Small Models as Experimental Subjects
Apart from questions towards the technical details of impossible languages, it is also important to explain the use of GPT-2 small models, instead of stronger models. We believe that GPT-2 small models play an essential role in language acquisition experiments of machine learning as guinea pigs or white mice in biological ones.
@@
In general, studying language learning on GPT-2 is analogous to experiments on white mouse to test vaccines. Both utilize the controllability in puny experimental subjects, and generalize findings from a miniature test to a larger scale. This operation represents the introduction of scientific and empirical methodology in AI.